The highest waterfall in the world is the Salto Angel Falls in Venezuela. Its longest fall has a height of 807 meters. If water at the top of the falls is at rest and at 15°C, what is the maximum temperature of the water at the bottom of the falls? Assume all the kinetic energy of the water as it reaches the bottom goes into raising its temperature. The specific heat capacity for water is 4186 J/kg°C.

3. The highest waterfall in the world is the Salto Angel Falls in Venezuela. Its longest fall has a height of 807 meters. If water at the top of the falls is at rest and at 15C, what is the maximum temperature of the water at the bottom of the falls? Assume all the kinetic energy of the water as it reaches the bottom goes into raising its temperature. The specific heat capacity for water is 4186 Jkg-C °
Answer #1

Mgh = change in kinetic energy

Change in kinetic energy = nn.Cv.dt

1×807×10 = 1×4186×(T-15)

T - 15 = 1.92

Final temperature = 16.92 degree Celsius